OpenStudy (anonymous):

How do I solve the lim of (secx - tanx) as x --> pi/2?

OpenStudy (noelgreco):

Change it to \[\lim _{x \rightarrow \frac{ \pi }{ 2 }}\frac{ 1 }{ \cos(x) }-\frac{ \sin(x) }{ \cos(x) }\] combine terms. Indeterminate as 0/0 -> apply l'Hopital.

OpenStudy (phi):

you could write sec x as 1/cos x, and tan x as sin x / cos x so \[ \sec x - \tan x = \frac{1- \sin x}{\cos x} \] multiply top and bottom by 1+ sin x what do you get ?

OpenStudy (anonymous):

(1-sin^2 x ) / (cosx)(1+sinx)

OpenStudy (anonymous):

then i have cosx / 1+sinx

OpenStudy (phi):

now take the limit

OpenStudy (anonymous):

the answer is 0?

OpenStudy (phi):

yes it approaches 0/2 = 0

OpenStudy (anonymous):

thank you so much

OpenStudy (phi):

yw